The equation describing a transverse wave on a string is
y(x,t)=( 2.50mm )sin[( 168s?1 )t?( 42.1m?1 )x].

A. Find the wavelength of this wave.

B. Find the frequency of this wave.

C. Find the amplitude of this wave.

D. Find the speed of motion of the wave.

E. Find the direction of motion of the wave.

F. Find the transverse displacement of a point on the string when t = 0.160s and at a position x = 0.140m.
Answer #1

comparing with y = Asin(wt - kx)

k = 2*pi/lambda= 42.1

wavelength = lambda = 2*3.14/42.1 = 0.15 m


w = 2*pi*f = 168


f = 26.75 Hz


A)


lambda = 0.15 m


B)f = 26.75 Hz


C) A = 2.5mm


D) v = w/k = (168)/42.1 = 3.99 m/s = 4 m/s


E) +x axis


F) y = 2.5mmsin((168*0.16)-(42.1*0.14)) = 2.11 mm